# thanks: Corresponding author

### v2

@@ Abstract
We investigate the pseudoscalar ($\eta_{t}$) and vector ($\psi_{t}$) toponium states, as well as the triply-top baryon ($\Omega_{ttt}$), using the QCD sum-rule method. This study was motivated by the recent observation of a pseudoscalar enhancement near the $t\bar{t}$ threshold, reported by the CMS and ATLAS collaborations with a statistical significance exceeding $5\sigma$. In the calculations, we consider both the perturbative and nonperturbative contributions, with the nonperturbative operators taken into account up to dimension eight. The results obtained for the pseudoscalar toponium align with the CMS prediction within the corresponding uncertainty range. The calculated negative binding energy for both the pseudoscalar and vector toponium states reflects the strong correlation within the $t\bar{t}$ system and can be interpreted as $t\bar{t}$ bound states. The results of this study can provide a precise theoretical guide for future experimental investigations of these states, which are composed entirely of top quarks, at high-energy colliders such as the LHC and future facilities like the FCC.
@@ I Introduction
Recent experimental findings have overturned the long-standing assumption that bound states of the top quark cannot form. In particular, the CMS CMS:2025kzt and ATLAS ATLAS:2025mvr Collaborations at the Large Hadron Collider (LHC) reported the first experimental evidence of a bound state of toponium. A key quantum signature accompanying the formation of this bound state, even at the highest accessible mass scales, is the observation of strong quantum entanglement between the spins of the top and anti-top quarks. This entanglement emphasizes that the $t\bar{t}$ system, despite the extremely short lifetime of its constituents, exists as a single, inseparable quantum entity. This discovery, achieved with a statistical significance exceeding $5\sigma$ CMS:2025kzt through the observation of its decay products, rapidly motivated a wide range of theoretical investigations Fuks:2025toq ; Matsuoka:2025jgm ; Sjostrand:2025qez ; Fuks:2025wtq ; Goncalves:2025hyx ; Zhu:2025ezg ; Zhang:2025fdp ; Lopez:2025kog ; Thompson:2025cgp ; Zhang:2025xxd ; Shao:2025dzw ; Bai:2025buy ; Ellis:2025nkm ; LeYaouanc:2025mpk ; Xiong:2025iwg ; Fu:2025yft ; Fu:2025zxb ; Fuks:2025sxu ; Afik:2025ejh ; Luo:2025psq . It is noteworthy that numerous theoretical predictions had previously anticipated the existence and properties of toponium Fadin:1987wz ; Kuhn:1987ty ; Barger:1987xg ; Strassler:1990nw ; Fadin:1990wx ; Fadin:1994pj ; Hoang:2000yr ; Hagiwara:2008df ; Penin:2005eu ; Sumino:2010bv ; Kiyo:2008bv ; Beneke:2015kwa ; Kawabata:2016aya ; Reuter:2018rbq ; Fuks:2021xje ; Aguilar-Saavedra:2024mnm ; Jiang:2024fyw ; Jafari:2025rmm ; Francener:2025tor ; dEnterria:2025ecx ; Garzelli:2024uhe ; Wang:2024hzd ; Fuks:2024yjj .
@@ III QCD Sum-Rule Mass Analysis
respectively. This mass deficit reflects the presence of strong quantum correlations within the $t\bar{t}$ system. In the quantum-field picture, these correlations may be interpreted as possible manifestations of spin–spin entanglement between the top and anti-top quarks, which could play a role in enhancing the dynamical stability of the bound state. Such effects, if confirmed, would represent a distinctive quantum feature of the toponium system. In contrast to the two-body toponium states, the triply-top baryon ($\Omega_{ttt}$) is a three-quark system that must satisfy the color-singlet constraint, stabilized by strong inter-quark interactions. While the calculated central mass, $m_{\Omega_{ttt}}=517.81{}^{+1.82}_{-1.88}\,\mathrm{GeV}$, slightly exceeds the sum of the constituent top-quark masses, $\sum m_{t}=517.68{}^{+0.93}_{-0.93}\,\mathrm{GeV}$, the lower bound of the uncertainty range falls below this threshold. This behavior should not be interpreted as a physically negative binding energy, but rather as a reflection of the intrinsic uncertainties of the QCD sum-rule framework, including variations of the Borel mass and continuum threshold parameters. From a qualitative standpoint, one may speculate that quantum correlations or multi-body entanglement among the three top quarks could play a role in maintaining the internal colorless configuration and stabilizing the baryonic state, although a more detailed theoretical and experimental investigation is required to clarify these effects. To provide a comprehensive comparison in light of recent measurements of the $t\bar{t}$ bound state, Table 3 summarizes the mass estimates for $\eta_{t}$, $\psi_{t}$, and $\Omega_{ttt}$ from various experimental and theoretical approaches, showing a notable consistency within the reported uncertainties. The overall consistency among different approaches provides additional confidence in the reliability of the present analysis.
@@
Table 3: Masses of the pseudoscalar ($\eta_{t}$) and vector ($\psi_{t}$) toponium, and the triply-top baryon ($\Omega_{ttt}$), compared with experimental results and theoretical predictions from various approaches.
@@ IV Conclusion
In this study, the masses of heavy hadrons containing the top quark—including the pseudoscalar toponium ($\eta_{t}$), vector toponium ($\psi_{t}$), and the triply-top baryon ($\Omega_{ttt}$)—have been investigated within the framework of QCD sum rules. The obtained mass results for $\eta_{t}$ and $\psi_{t}$ show good agreement with recent experimental measurements and are consistent with existing theoretical predictions and phenomenological estimates. Another notable outcome of this work is the prediction of the mass of the triply-top baryon $\Omega_{ttt}$, representing the first precise QCD sum-rule determination for this ultraheavy system. Furthermore, our numerical analysis of the OPE indicates a characteristic pattern in the relative contributions of different terms in heavy-quark systems. In particular, the dimension $D=4$ nonperturbative term dominates in the baryonic channel $\Omega_{ttt}$, while in the toponium channels a similar dominance is observed for the dimension $D=6$ term. This observation highlights the notable role of nonperturbative effects at very high quark-mass scales.

### v3

@@ Abstract
We investigate the pseudoscalar ($\eta_{t}$) and vector ($\psi_{t}$) toponium states, as well as the triply-top baryon ($\Omega_{ttt}$), using the QCD sum-rule method. This study was motivated by the recent observation of a pseudoscalar enhancement near the $t\bar{t}$ threshold, reported by the CMS and ATLAS collaborations with a statistical significance exceeding $5\sigma$. In the calculations, we consider both the perturbative and nonperturbative contributions, with the nonperturbative operators taken into account up to dimension eight. The results obtained for the pseudoscalar toponium provide a theoretical estimate that is consistent with the near-threshold events observed in recent experimental studies. The calculated negative binding energy for both the pseudoscalar and vector toponium states reflects the strong correlation within the $t\bar{t}$ system and can be interpreted as $t\bar{t}$ bound states, while the calculated central mass for the $\Omega_{ttt}$ slightly exceeds the central value of the sum of the constituent top-quark masses. The results of this study can provide a precise theoretical guide for future experimental investigations of these states, which are composed entirely of top quarks, at high-energy colliders such as the LHC and future facilities like the FCC.
@@ I Introduction
Recent experimental findings have called into question the long-standing assumption that bound states of the top quark cannot form. In particular, the CMS CMS:2025kzt and ATLAS ATLAS:2025mvr Collaborations at the LHC reported a statistically significant excess near the $t\bar{t}$ threshold, consistent with the production of a pseudoscalar quasi-bound toponium state. A key quantum signature of top–anti-top systems near threshold, even at the highest accessible mass scales, is the strong spin correlation between the top and anti-top quarks, which may reflect quasi-bound dynamics or significant quantum entanglement. Such entanglement, although not directly observed, may influence the behavior of the $t\bar{t}$ system near threshold, particularly given the extremely short lifetimes of its constituents. This experimental observation, achieved with a statistical significance exceeding $5\sigma$ CMS:2025kzt through the observation of its decay products, rapidly motivated a wide range of theoretical investigations Fuks:2025toq ; Matsuoka:2025jgm ; Sjostrand:2025qez ; Fuks:2025wtq ; Goncalves:2025hyx ; Zhu:2025ezg ; Zhang:2025fdp ; Lopez:2025kog ; Thompson:2025cgp ; Zhang:2025xxd ; Shao:2025dzw ; Bai:2025buy ; Ellis:2025nkm ; LeYaouanc:2025mpk ; Xiong:2025iwg ; Fu:2025yft ; Fu:2025zxb ; Fuks:2025sxu ; Afik:2025ejh ; Luo:2025psq . It is noteworthy that numerous theoretical predictions had previously anticipated the existence and properties of toponium Fadin:1987wz ; Kuhn:1987ty ; Barger:1987xg ; Strassler:1990nw ; Fadin:1990wx ; Fadin:1994pj ; Hoang:2000yr ; Hagiwara:2008df ; Penin:2005eu ; Sumino:2010bv ; Kiyo:2008bv ; Beneke:2015kwa ; Kawabata:2016aya ; Reuter:2018rbq ; Fuks:2021xje ; Aguilar-Saavedra:2024mnm ; Jiang:2024fyw ; Jafari:2025rmm ; Francener:2025tor ; dEnterria:2025ecx ; Garzelli:2024uhe ; Wang:2024hzd ; Fuks:2024yjj .
@@ III QCD Sum-Rule Mass Analysis
respectively. This mass deficit reflects the presence of strong quantum correlations within the $t\bar{t}$ system. In the quantum-field picture, these correlations may be interpreted as possible manifestations of spin–spin entanglement between the top and anti-top quarks, which could play a role in enhancing the dynamical stability of the bound state. Such effects, if confirmed, would represent a distinctive quantum feature of the toponium system. In contrast to the two-body toponium states, the triply-top baryon ($\Omega_{ttt}$) is a three-quark system that must satisfy the color-singlet constraint, stabilized by strong inter-quark interactions. While the calculated central mass, $m_{\Omega_{ttt}}=517.81{}^{+1.82}_{-1.88}\,\mathrm{GeV}$, slightly exceeds the sum of the constituent top-quark masses, $\sum m_{t}=517.68{}^{+0.93}_{-0.93}\,\mathrm{GeV}$, the lower bound of the uncertainty range falls below this threshold. This behavior should not be interpreted as a physically negative binding energy, but rather as a reflection of the intrinsic uncertainties of the QCD sum-rule framework, including variations of the Borel mass and continuum threshold parameters. From a qualitative standpoint, one may speculate that quantum correlations or multi-body entanglement among the three top quarks could play a role in maintaining the internal colorless configuration and stabilizing the baryonic state, although a more detailed theoretical and experimental investigation is required to clarify these effects. The CMS collaboration has recently reported an excess of events near the $t\bar{t}$ threshold in final states with two charged leptons and multiple jets, using proton–proton collision data at $\sqrt{s}=13~\mathrm{TeV}$ corresponding to an integrated luminosity of 138 fb-1 CMS:2025kzt . This excess, observed with a statistical significance exceeding $5\sigma$, was interpreted as evidence for a quasi-bound pseudoscalar toponium state ($^{1}\mathrm{S}_{0}^{[1]}$}). The measured cross section relative to the fixed-order pQCD prediction is $8.8^{+1.2}_{-1.4}$ pb. No direct mass measurement of the pseudoscalar toponium was provided, though these results offer valuable guidance for theoretical studies. In parallel with the CMS results, the ATLAS Collaboration performed an independent and complementary analysis of $t\bar{t}$ production near the kinematic threshold using 140 fb-1 of proton–proton collision data at $\sqrt{s}=13$ TeV. The study focused on dileptonic final states with multiple jets and compared the observed invariant-mass spectrum to both a baseline perturbative QCD prediction and an extended model incorporating simulations of colour-singlet quasi-bound state formation. A statistically significant excess of events was observed over the baseline expectation, with a significance of 7.7 standard deviations, consistent with the production of an S-wave colour-singlet quasi-bound $t\bar{t}$ state as predicted by non-relativistic QCD. The corresponding excess cross section was measured to be $9.0\pm 1.3$ pb, further supporting the interpretation of near-threshold dynamics beyond pure perturbative descriptions. To facilitate a comprehensive comparison in light of recent $t\bar{t}$ measurements, Table 3 presents the mass estimates for $\eta_{t}$, $\psi_{t}$, and $\Omega_{ttt}$ from various theoretical studies, demonstrating notable consistency within the reported uncertainties. This overall agreement further reinforces the reliability of the present analysis.
@@
Table 3: Masses of the pseudoscalar ($\eta_{t}$) and vector ($\psi_{t}$) toponium, and the triply-top baryon ($\Omega_{ttt}$), compared with theoretical predictions from various approaches.
@@ IV Conclusion
In this study, the masses of heavy hadrons containing the top quark—including the pseudoscalar toponium ($\eta_{t}$), vector toponium ($\psi_{t}$), and the triply-top baryon ($\Omega_{ttt}$)—have been investigated within the framework of QCD sum rules. The obtained mass results for $\eta_{t}$ and $\psi_{t}$ are consistent with existing theoretical predictions and phenomenological estimates, and in the case of $\eta_{s}$ aligns well with the constraints suggested by recent experimental observations. Another notable outcome of this work is the prediction of the mass of the triply-top baryon $\Omega_{ttt}$, representing the first precise QCD sum-rule determination for this ultraheavy system. Furthermore, our numerical analysis of the OPE indicates a characteristic pattern in the relative contributions of different terms in heavy-quark systems. In particular, the dimension $D=4$ nonperturbative term dominates in the baryonic channel $\Omega_{ttt}$, while in the toponium channels a similar dominance is observed for the dimension $D=6$ term. This observation highlights the notable role of nonperturbative effects at very high quark-mass scales.